Circular No. 8252

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003kw IN UGC 6314 Further to IAUC 8245, M. Armstrong reports his discovery of an ap-

parent supernova (mag 18.5) on unfiltered CCD images taken with a 0.35-m reflector on Dec. 7.143 and 8.206 UT. The new object is located at α = 11^{h}18^{m}12^{s}.21, δ = +30^{o}23′41′′.5 (equinox 2000.0), which is ≈ 7′′.8 east and 1′′.5 south of the center of UGC 6314. Nothing was visible on Armstrong’s image from Apr. 1 (limiting mag 19.5) or Palomar Sky Survey red and blue plates (limiting mag 20.5).

2003 EH_{1} AND THE QUADRANTIDS P. Jenniskens, NASA Ames Research Center, has pointed out that 2003

EH_{1} (cf. MPEC 2003-E27) would seem to be a very strong candidate for the parent of the Quadrantid meteor stream. The later orbits, from arcs of up to 48 days (MPO 48330), indicate that frequent approaches within 0.2–0.3 AU of Jupiter occur, those during the past century or two evidently increasing q from just under 1 AU (with other orbital elements also very similar to those of the Quadrantids) to the present 1.19 AU. The current theoretical radiant for 2003 EH_{1} (α = 229^{o}.9, δ = +49^{o}.6; V_{inf} = 41.7 km/s at solar longitude 282^{o}.94, equinox 2000.0) is at the center of the Quadrantid radiants measured by photographic means, the narrow dispersion implying a young (∼ 500 years) shower age. From that dispersion, Jenniskens et al. (1997, Astron. Astrophys. 327, 1242) suspected that the parent was still among the meteoroids, hiding as a minor planet. On computing a parabolic orbit for C/1490 Y1, Hasegawa (1979, Publ. Astron. Soc. Japan 31, 257) introduced that comet as the likely Quadrantid parent. In attempting to link the 2003 observations to those of 1490–1491, Jenniskens, and also B. G. Marsden (Center for Astrophysics), have found that most of the potential solutions with the required Jan. 1491 perihelion date yield 0.5 < q < 0.6 AU in 1491, and this is probably too small to fit the data used by Hasegawa. Values in the more acceptable range of 0.7 < q < 0.8 AU (and 0.80 > e > 0.75) certainly arise for 1488 < T < 1494, however, the desired date being clearly attainable with the help also of a close approach to the earth or — more likely — the presence of nongravitational forces. Further light could be shed on the problem by the recognition of precovery and/or recovery observations of 2003 EH_{1}, which is presumably a comet and that should in any case be considered a high-priority object for further study.

2003 December 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green